# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v4

@@ 3 HydroPML: Process-based Hydrology in Physics-aware Machine Learning
Figure 5: The proposed process-based hydrology in physics-aware machine learning (HydroPML). Red labels denote different objectives of HydroPML. For example, the oval labeled 1.1 represents short-term forecasts. See the text for more detail.
@@ 3.1 Rainfall-runoff Process Understanding
Rainfall-runoff models are extensively utilized in hydrology to investigate hydrological processes and play a crucial role in water resources management, encompassing runoff prediction [203], flood prediction [204], and drought analysis [205]. These models have been developed for diverse applications, ranging from small catchments to global scales. The hydrological model can be classified into three broad groups: the empirical model, conceptual models, and physical-based models [2]. Detailed information regarding Physical Methods in Hydrological Process is available in Appendix A of Supplementary Material.
@@ 3.1.1 Rainfall-runoff Forecast Meets Physics-aware Machine Learning
Table 5: PaML-based Rainfall-runoff Hydrological Forecasts Classified by Objectives
@@
(1.5) Missing Data and Ungauged Basins Attention-based deep learning (such as 3D-CNN-Transformer) [243], input-selection ensemble method [244] AI-based hybrid models [196], serial hybrid LSTMs [245, 246], transfer learning in SWAT-LSTM [217].
@@
Short-term Forecasts. See Fig. 5 (label 1.1). Short-term forecasts focus on outlook horizons ranging from hours to weeks. Innovative forecasting tools are transforming flood early warning mechanisms, and agricultural and hydropower management schemes. PaML offers robust tools to address the formidable challenges associated with the uncertainties, data dependencies, and intricacies inherent in short-term hydrological forecasting. First, the significant uncertainty and complexity inherent in hydrological processes and weather-climate factors affecting river basins have increasingly motivated researchers to adopt PDgML approaches. Specifically, the PDgML models based on LSTM [203, 206] are proposed to predict streamflow. They have been found to have greater accuracy in predicting hourly and daily streamflow than a classical neural network with back-propagation. A model [207] utilizing multiple GRUs is proposed for predicting streamflow up to 120 hours ahead. Granata et al. [211] demonstrate that their ensemble model, based on random forest and multilayer perceptron, outperforms bi-directional LSTM networks in predicting peak flow rates, while also achieving significantly shorter computation times. Wegayehu and Muluneh [210] compare various ML algorithms for one-step daily streamflow forecasting, showing that both MLP and GRU models perform better than stacked LSTM and bi-directional LSTM models to a similar extent. An ensemble model [212] incorporating nonlinear autoregressive networks, multilayer perceptrons, and random forests is proposed for short-term streamflow forecasting, using precipitation as the sole exogenous input, with a forecast horizon of up to 7 days. In addition, to model the spatial diversity of rivers, the graph convolutional GRUs-based model [208] is designed to predict streamflow for the next 36 hours at a sensor location by utilizing data from the upstream river network. Graph convolutional RNN [209] is developed for robust water demand forecasting. More detailed information about PDgML studies on short-term streamflow prediction can be found in [261]. Furthermore, to overcome data dependence, attempts have been made based on the PiML community, such as PINNs, for solving groundwater flow equations [215]. Because it is difficult to express the whole process of rainfall-runoff with specific equations, the current progress based on PiML technology is very slow. In addition, integrating hydrologic processes with LSTM [216] proves effective in assimilating recent streamflow observations, thereby enhancing near-term daily streamflow forecasts. A coupled soil and water assessment tool (SWAT)-LSTM approach [217] is developed to provide a potential shortcut for conducting daily streamflow simulations in both ungauged and poorly gauged watersheds. The conceptual hydrological model by integrating NNs [218] reveals that the hybrid approach outperforms both the original GR4J model and the single NN-based runoff prediction model in terms of prediction accuracy.
@@
Spatial Variability. See Fig. 5 (label 1.3). Rainfall-runoff modeling is difficult due to the non-linear structure of the process, which involves much spatio-temporal variability. Spatial variability is critical for predictions in ungauged basins and for developing universal regional models. PaML (especially GNNs) provides the possibility of distributed modeling. Specifically, for PDgML, GNN [228] is used for fully-distributed rainfall-runoff modeling. The result shows that the GNN-based fully-distributed model [228] successfully represents the spatial information in predictions. A graph convolutional GRUs based model [208], ConvLSTM [213], and graph convolutional RNN [209] can also be developed to model spatial diversity of rivers. A recurrent GNN [231] is proposed to capture the interactions among multiple segments in the river network. In this study, a physics-informed model is employed to initialize the ML model and capture the underlying physics of streamflow. To mitigate the requirement for large training datasets, a real-time active learning approach is utilized within a graph-based reinforcement learning framework [229]. This method leverages spatial and temporal contextual information to select representative query samples. In addition, region-specific differences can be learned for the LSTM model by inputting sufficient heterogeneous data (data synergy) [230].
@@
Bias Reduction and Reliability. See Fig. 5 (label 1.4). One important challenge of PaML-based rainfall-runoff forecasts is the ability of such models to produce reliable or physically-plausible results in different domains, especially extreme conditions such as floods and droughts. Specifically, reliable spatio-temporal remote sensing data and directly measured datasets can be used to reduce the bias of PDgML-based rainfall-runoff forecasting. For example, bias-corrected remote-sensing precipitation products are adopted as precipitation input of a spatio-temporal DL rainfall-runoff forecasting model [235], which can improve the model’s reliability for extreme flood forecasts. Directly measured datasets are utilized for streamflow estimation to develop a data-driven forecasting algorithm [214]. However, these bias-corrected data-based methods face challenges over small catchments due to uncertainties in remote sensing data and errors in hydrological models. Thus, some ensemble methods (such as resampling ensemble methods [236], which combine hydrometeorological modeling and LSTM [237]) are explored to improve the reliability and robustness of PDgML models. Furthermore, one emerging route for bias reduction and reliability is to employ PeML designs that explicitly observe the physical laws in the rainfall-runoff process, such as MC-LSTM [168] and differentiable modeling [13]. Another new development involves hybrid models designed to generate physically plausible or explainable results. For instance, Wi et al. [239] explore the capability of DL models to provide reliable future projections of streamflow under warming conditions. It concludes that integrating estimates of evapotranspiration from process-based models as additional input features can enhance the performance of LSTM models in hydrologic projections under climate change. Alternatively, training LSTMs on a diverse set of watersheds also proves effective. Another approach involves coupling physical models (such as global climate models or global hydrological models) with LSTM models [240] for long-term streamflow projections. Here, the LSTM acts as a post-processor aimed at constraining streamflow simulations derived from physics-based models to reduce uncertainty. To improve streamflow prediction within the Weather Research and Forecasting Hydrological Modeling System (WRF-Hydro), a method known as WRF-Hydro-LSTM [241] is proposed. This method utilizes LSTM models to predict residual errors within WRF-Hydro’s simulations.
@@
However, due to data errors in hydrological records, an unconstrained ML (such as LSTM) outperforms a PeML model (such as MC-LSTM) because of the ability of PDgML to learn and accommodate these data errors [238]. In addition, three LSTM daily streamflow prediction models [242] are built and compared, including a LSTM post-processor trained on outputs from the United States National Water Model (NWM), a LSTM post-processor trained on NWM outputs and atmospheric forcings, and a LSTM model trained solely on atmospheric forcings. It concludes that a LSTM model trained only on atmospheric forcing outperforms the other two LSTMs in ungauged basins. Hybrid models may reduce the reliability of the model in some cases. Thus, PiML, PeML, and PaHL approaches for the reliability of rainfall-runoff forecasts require further development in the future.
@@
Missing Data and Ungauged Basins. See Fig. 5 (label 1.5). Hydrologic predictions at many watersheds, such as rural watersheds, as well as cold and ungauged regions, are important but also challenging due to data scarcity [262]. Thus, many PaML-based methods are developed for poorly gauged or ungauged basins. For example, geo-spatio-temporal mesoscale data and attention-based deep learning [243] are used to improve long-term streamflow prediction. This study demonstrates the superior performance of the 3D-CNN–Transformer compared to TD-CNN–Transformer, TD-CNN–LSTNet, and 3D-CNN–LSTNet in hydroclimate data applications, particularly for poorly gauged basins. Despite the excellent performance of deep neural networks in streamflow predictions, their effectiveness declines in ungauged regions. To mitigate these errors, an input-selection ensemble method [244] is proposed, leveraging the flexibility of deep networks to integrate satellite-based soil moisture products or daily flow distributions. Furthermore, hybrid models present viable alternatives to hydrological models, especially in watersheds lacking measured data such as climatic parameters, land cover, soil data, and groundwater aquifer properties, provided that appropriate inputs like rainfall are available. Mohammadi et al. [196] demonstrate that AI-based hybrid models generally yield more accurate streamflow estimates than conceptual rainfall-runoff models in ungauged basins. To improve out-of-distribution prediction in data-scarce basins, several sequential methods have been proposed for developing hybrid LSTMs [245, 246]. These methods involve using outputs from physics-based hydrologic model simulations as additional inputs in LSTM networks. Additionally, to model long-duration daily streamflow in poorly gauged watersheds, transfer learning techniques are employed with a pre-trained SWAT-LSTM model [217].
@@
However, these PaML-based approaches are still only preliminary attempts. Deeply exploiting the transferability of ML models (such as domain adaptation methods [263], meta-learning methods [264]) and the interpretability of physical models will effectively solve hydrological prediction problems in ungauged basins. Furthermore, leveraging the increasing volume of simulation and observation data in hydrology, we can integrate large pre-trained models (e.g., generative models in ChatGPT [265], Transformer in ClimaX [266]) with hydrological knowledge to establish a foundational hydrology pre-training model. The pre-training large model can then be finely tuned to address diverse hydrological prediction tasks, including those in ungauged basins.
@@
Parameterization. See Fig. 5 (label 1.6). Parameters for predicting rainfall-runoff are typically not directly measurable; instead, they are indirectly estimated through prior knowledge or model calibration. Calibration involves adjusting model parameters to match outputs with observed data at specific locations. Specifically, a ML approach and remotely sensed underlying surface data [247] are employed to develop models for estimating the runoff routing parameter. This study demonstrates that ML-derived parameters from extensive datasets exhibit robustness, ensuring the overall performance of physical models. A deep neural network [248] is used to construct inverse mapping from informative responses to each of the selected parameters. Yang et al. [249] calibrate three essential parameters of the Variable Infiltration Capacity model at every 1/8°$\degree$ grid-cell employing ML-based maps (shuffled complex evolution algorithm). Furthermore, some PaHL-based methods, such as a differentiable programming framework [202] and differentiable parameter learning [227], are proposed to efficiently learn a mapping between inputs and parameters of process-based models in an end-to-end manner.
@@
Uncertainty Estimation in Rainfall-runoff Forecasts. See Fig. 5 (label 1.7). Hydrological forecasts often suffer from inaccuracies due to insufficient conceptualization of underlying physics, inaccurate predictions of ML models, non-uniqueness of model parameters, and uncertainties in calibration data, such as streamflow outputs. Therefore, precise uncertainty estimations are crucial for actionable hydrological predictions. However, the majority of PaML-based rainfall-runoff studies do not provide uncertainty estimates. In order to address this problem, some PDgML approaches are proposed for model uncertainty, such as a data-driven approach [250], Bayesian processor [251], a hybrid ensemble and variational data assimilation framework [252], a hierarchical model tree [253], and probabilistic deep ensembles [254]. In addition, several LSTM-based approaches are presented for uncertainty estimation. Two methods [255] are explored for integrating LSTM with Gaussian processes. In the first method, LSTM parameterizes a Gaussian process, while in the second, LSTM acts as a forecast model with a Gaussian process post-processor. Althoff et al. [althoff2022predictive] investigate predictive uncertainty in hydrological models based on LSTM by comparing multiparameter ensembles with dropout ensembles. Song et al. [song2023framework] propose a framework or quantifying uncertainty contributions from the sample set, ML method, ML architecture, and their interactions in multi-step time-series forecasting using variance analysis. Klotz et al. [klotz2023uncertainty] establish an uncertainty estimation benchmarking procedure and present four PDgML (LSTM) baselines, including three based on mixture density networks and one on Monte Carlo dropout. In addition, by integrating physics and ML approaches, PaHL has also been gaining attention, as it can alleviate the computational burden and improve efficiency for uncertainty estimation. For example, the LSTM model is considered a post-processor [240] that aims to reduce the uncertainty of streamflow simulations from the physics-based model. Serial hybrid LSTM [246] is introduced for uncertainty quantification. An integrated model (XAJ-MCQRNN) [259] is developed, incorporating the Xinanjiang conceptual model and composit e quantile regression neural network, to address error propagation and accumulation in multi-step flood probability density forecasts. Additionally, a hybrid error correction model [260] is proposed to enhance streamflow forecast accuracy. The hybrid model integrates the HBV model with the ML algorithm through a data assimilation technique.
In general, integrating physics knowledge into ML for uncertainty estimation has the potential to better characterize uncertainty. For instance, ML surrogate models might yield physically inconsistent predictions; integrating physical principles can help alleviate this problem. Moreover, reducing data requirements of ML models due to constraints imposed by known rainfall-runoff processes could alleviate some of the computational costs associated with NNs.

### v5

@@ 3 HydroPML: Process-based Hydrology in Physics-aware Machine Learning
Figure 5: The proposed process-based hydrology in physics-aware machine learning (HydroPML). HydroPML encompasses rainfall-runoff hydrological process understanding over time scales ranging from hours to decades and hydrodynamic process understanding over time scales from seconds to days. Red labels denote different objectives of HydroPML. For example, the oval labeled 1.1 represents short-term forecasts. The lower left and right corners represent the connections between PaML and the objects for rainfall-runoff hydrological processes and hydrodynamic processes, respectively. See the text for more detail.
@@ 3.1 Rainfall-runoff Process Understanding
Rainfall-runoff models are extensively utilized in hydrology to investigate hydrological processes and play a crucial role in water resources management, encompassing runoff prediction [206], flood prediction [207], and drought analysis [208]. These models have been developed for diverse applications, ranging from small catchments to global scales. The hydrological model can be classified into three broad groups: the empirical model, conceptual models, and physical-based models [2]. Detailed information regarding Physical Methods in Hydrological Process is available in Appendix A of SM.
@@ 3.1.1 Rainfall-runoff Forecast Meets Physics-aware Machine Learning
Table 2: PaML-based rainfall-runoff hydrological forecasts classified by objectives.
@@
(1.5) Missing Data and Ungauged Basins Attention-based deep learning (such as 3D-CNN-Transformer) [246], input-selection ensemble method [247], encoder–decoder LSTMs [248], encoder-decoder double-layer LSTM [249] AI-based hybrid models [198], serial hybrid LSTMs [250, 251], transfer learning in SWAT-LSTM [220].
@@
Short-term Forecasts. See Fig. 5 (label 1.1). Short-term forecasts focus on outlook horizons ranging from hours to weeks. Innovative forecasting tools are transforming flood early warning mechanisms, and agricultural and hydropower management schemes. PaML offers robust tools to address the formidable challenges associated with the uncertainties, data dependencies, and intricacies inherent in short-term hydrological forecasting. First, the significant uncertainty and complexity inherent in hydrological processes and weather-climate factors affecting river basins have increasingly motivated researchers to adopt PDgML approaches. Specifically, the PDgML models based on LSTM [206, 209] are proposed to predict streamflow. They have been found to have greater accuracy in predicting hourly and daily streamflow than a classical NN with back-propagation. A model [210] utilizing multiple GRUs is proposed for predicting streamflow up to 120 hours ahead. Granata et al. [214] demonstrate that their ensemble model, based on random forest and multilayer perceptron, outperforms bi-directional LSTM networks in predicting peak flow rates, while also achieving significantly shorter computation times. Wegayehu and Muluneh [213] compare various ML algorithms for one-step daily streamflow forecasting, showing that both MLP and GRU models perform better than stacked LSTM and bi-directional LSTM models. An ensemble model [215] incorporating nonlinear autoregressive networks, multilayer perceptrons, and random forests is proposed for short-term streamflow forecasting, using precipitation as the only exogenous input, with a forecast horizon of up to 7 days. In addition, to model the spatial diversity of rivers, the graph convolutional GRUs-based model [211] is designed to predict streamflow for the next 36 hours at a sensor location by utilizing data from the upstream river network. Graph convolutional RNN [212] is developed for robust water demand forecasting. More detailed information about PDgML studies on short-term streamflow prediction can be found in [266]. Furthermore, to overcome data dependence, attempts have been made based on the PiML community, such as PINNs, for solving groundwater flow equations [218]. Because it is difficult to express the whole process of rainfall-runoff with specific equations, the current progress based on PiML technology is very slow. In addition, integrating hydrologic processes with LSTM [219] proves effective in assimilating recent streamflow observations, thereby enhancing near-term daily streamflow forecasts. A coupled soil and water assessment tool (SWAT)-LSTM approach [220] is developed to provide a potential shortcut for conducting daily streamflow simulations in both ungauged and poorly gauged watersheds. The conceptual hydrological model by integrating NNs [221] reveals that the hybrid approach outperforms both the original GR4J model and the single NN-based runoff prediction model in terms of prediction accuracy.
@@
Spatial Variability. See Fig. 5 (label 1.3). Rainfall-runoff modeling is difficult due to the non-linear structure of the process, which involves much spatio-temporal variability. Spatial variability is critical for predictions in ungauged basins and for developing universal regional models. PaML (especially GNNs) provides the possibility of distributed modeling. Specifically, for PDgML, GNN [231] is used for fully-distributed rainfall-runoff modeling. The result shows that the GNN-based fully-distributed model [231] successfully represents the spatial information in predictions. A graph convolutional GRUs based model [211], ConvLSTM [216], and graph convolutional RNN [212] can also be developed to model spatial diversity of rivers. A recurrent GNN [234] is proposed to capture the interactions among multiple segments in the river network. In this work, a physics-based model is used to initialize the ML model and learn the physics of streamflow. To mitigate the requirement for large training datasets, a real-time active learning approach is employed in a graph-based reinforcement learning framework [232]. This method leverages spatial and temporal contextual information to select representative query samples. In addition, region-specific differences can be learned for the LSTM model by inputting sufficient heterogeneous data (data synergy) [233].
@@
Bias Reduction and Reliability. See Fig. 5 (label 1.4). One important challenge of PaML-based rainfall-runoff forecasts is the ability of such models to produce reliable or physically-plausible results in different domains, especially extreme conditions such as floods and droughts. Specifically, reliable spatio-temporal remote sensing data and directly measured datasets can be used to reduce the bias of PDgML-based rainfall-runoff forecasting. For example, bias-corrected remote-sensing precipitation products are adopted as precipitation input of a spatio-temporal DL rainfall-runoff forecasting model [238], which can improve the model’s reliability for extreme flood forecasts. Directly measured datasets are utilized for streamflow estimation to develop a data-driven forecasting algorithm [217]. However, these bias-corrected data-based approaches suffer from deficiencies over small catchments due to uncertainty in remote sensing data and errors in hydrological models. Thus, some ensemble methods (such as resampling ensemble methods [239], which combine hydrometeorological modeling and LSTM [240]) are explored to improve the reliability and robustness of PDgML models. Furthermore, one emerging route for bias reduction and reliability is to employ PeML designs that explicitly observe the physical laws in the rainfall-runoff process, such as MC-LSTM [166] and differentiable modeling [13]. Another new development is hybrid models to produce physically-plausible or explainable results. For instance, Wi et al. [242] explore the capability of DL models to provide reliable future projections of streamflow under warming conditions. It concludes that integrating estimates of evapotranspiration from process-based models as additional input features can enhance the performance of LSTM models in hydrologic projections under climate change. Alternatively, training LSTMs on a diverse set of watersheds also proves effective. Another approach involves coupling physical models (such as global climate models or global hydrological models) with LSTM models [243] for long-term streamflow projections. Here, the LSTM acts as a post-processor aimed at constraining streamflow simulations derived from physics-based models to reduce uncertainty. To improve streamflow prediction within the Weather Research and Forecasting Hydrological Modeling System (WRF-Hydro), a method known as WRF-Hydro-LSTM [244] is proposed. This method utilizes LSTM models to predict residual errors within WRF-Hydro’s simulations.
@@
However, due to data errors in hydrological records, an unconstrained ML (such as LSTM) outperforms a PeML model (such as MC-LSTM) because of the ability of PDgML to learn and accommodate these data errors [241]. In addition, three LSTM-based daily streamflow prediction models [245] are built and compared, including a LSTM post-processor trained on outputs from the United States National Water Model (NWM), a LSTM post-processor trained on NWM outputs and atmospheric forcings, and a LSTM model trained solely on atmospheric forcings. It concludes that a LSTM model trained only on atmospheric forcing outperforms the other two LSTMs in ungauged basins. Hybrid models may reduce the reliability of the model in some cases. Thus, PiML, PeML, and PaHL approaches for the reliability of rainfall-runoff forecasts require further development in the future.
@@
Missing Data and Ungauged Basins. See Fig. 5 (label 1.5). Hydrologic predictions at many watersheds, such as rural watersheds, as well as cold and ungauged regions, are important but also challenging due to data scarcity [267]. Thus, many PaML-based methods are developed for poorly gauged or ungauged basins. For example, geo-spatio-temporal mesoscale data and attention-based deep learning [246] are used to improve long-term streamflow prediction. This study demonstrates the superior performance of the 3D-CNN–Transformer compared to TD-CNN–Transformer, TD-CNN–LSTNet, and 3D-CNN–LSTNet in hydroclimate data applications, particularly for poorly gauged basins. Despite the excellent performance of deep NNs in streamflow predictions, their effectiveness declines in ungauged regions. To mitigate these errors, an input-selection ensemble method [247] is proposed, leveraging the flexibility of deep networks to integrate satellite-based soil moisture products or daily flow distributions. Recently, a watershed-aware streamflow forecast model, an encoder-decoder double-layer LSTM [249], is proposed by encoding basin spatial properties into LSTM. This model demonstrates its effectiveness in data-scarce regions such as Chile. Concurrently, another model, encoder–decoder LSTMs [248], is developed for streamflow forecasts at a global scale. This encoder–decoder model reliably predicts extreme river events with a lead time of up to 5 days in ungauged basins. These advancements highlight the significant potential of PDgML for streamflow and flood forecasting in ungauged basins. Furthermore, hybrid models present viable alternatives to hydrological models, especially in watersheds lacking measured data such as climatic parameters, land cover, soil data, and groundwater aquifer properties, provided that appropriate inputs like rainfall are available. Mohammadi et al. [198] demonstrate that AI-based hybrid models generally yield more accurate streamflow estimates than conceptual rainfall-runoff models in ungauged basins. Additionally, the hybrid framework [268] combines a conceptual hydrological model (such as Glacial Snow Melt (GSM)) with Support Vector Regression (SVR) and firefly algorithm-driven parameter optimization, demonstrating superior performance in rainfall runoff prediction compared to standalone GSM and conventional SVR. To improve out-of-distribution prediction in data-scarce basins, several sequential methods have been proposed for developing hybrid LSTMs [250, 251]. These methods involve using outputs from physics-based hydrologic model simulations as additional inputs in LSTM networks. Additionally, to model long-duration daily streamflow in poorly gauged watersheds, transfer learning techniques are employed with a pre-trained SWAT-LSTM model [220].
@@
However, these PaML-based approaches are still only preliminary attempts. Deeply exploiting the transferability of ML models (such as domain adaptation methods [269, 270], meta-learning methods [271]) and the interpretability of physical models will effectively solve hydrological prediction problems in ungauged basins. Furthermore, leveraging the increasing volume of simulation and observation data in hydrology, we can integrate large pre-trained models (e.g., generative models in ChatGPT [272], Transformer in ClimaX [273]) with hydrological knowledge to establish a foundational hydrology pre-training model. The pre-training large model can then be finely tuned to address diverse hydrological prediction tasks, including those in ungauged basins.
@@
Parameterization. See Fig. 5 (label 1.6). Parameters for predicting rainfall-runoff are typically not directly measurable; instead, they are indirectly estimated through prior knowledge or model calibration. Calibration involves adjusting model parameters to match outputs with observed data at specific locations. Specifically, a ML approach and remotely sensed underlying surface data [252] are employed to develop models for estimating the runoff routing parameter. This study demonstrates that ML-derived parameters from extensive datasets exhibit robustness, ensuring the overall performance of physical models. A deep neural network [253] is used to construct inverse mapping from informative responses to each of the selected parameters. Yang et al. [254] calibrate three essential parameters of the Variable Infiltration Capacity model at every 1/8 grid-cell employing ML-based maps (shuffled complex evolution algorithm). Furthermore, some PaHL-based methods, such as a differentiable programming framework [204] and differentiable parameter learning [230], are proposed to efficiently learn a mapping between inputs and parameters of process-based models in an end-to-end manner.
@@
Uncertainty Estimation in Rainfall-runoff Forecasts. See Fig. 5 (label 1.7). Hydrological forecasts often suffer from inaccuracies due to insufficient conceptualization of underlying physics, inaccurate predictions of ML models, non-uniqueness of model parameters, and uncertainties in calibration data (such as streamflow outputs). Therefore, precise uncertainty estimations are crucial for actionable hydrological predictions. However, the majority of PaML-based rainfall-runoff studies do not provide uncertainty estimates. In order to address this problem, some PDgML approaches are proposed for model uncertainty, such as a data-driven approach [255], Bayesian processor [256], a hybrid ensemble and variational data assimilation framework [257], a hierarchical model tree [258], and probabilistic deep ensembles [259]. In addition, several LSTM-based approaches are presented for uncertainty estimation. Two methods [260] are explored for integrating LSTM with Gaussian processes. In the first method, the LSTM is used to parameterize a Gaussian process, and in the second, the LSTM acts as a forecast model with a Gaussian process post-processor. Althoff et al. [261] investigate predictive uncertainty in hydrological models based on LSTM by comparing multiparameter ensembles with dropout ensembles. Song et al. [263] propose a framework or quantifying uncertainty contributions from the sample set, ML method, ML architecture, and their interactions in multi-step time-series forecasting using variance analysis. Furthermore, to reduce the gap between the hydrological sciences and ML community for uncertainty estimations, Klotz et al. [262] establish an uncertainty estimation benchmarking procedure and present four PDgML (LSTM) baselines, including three based on mixture density networks and one based on Monte Carlo dropout. In addition, by integrating physics and ML approaches, PaHL has also been gaining attention, as it can alleviate the computational burden and improve efficiency for uncertainty estimation. For example, the LSTM model is considered a post-processor [243] that aims to reduce the uncertainty of streamflow simulations from the physics-based model. Serial hybrid LSTM [251] is introduced for uncertainty quantification. An integrated model (XAJ-MCQRNN) [264] is developed, incorporating the Xinanjiang conceptual model and composite quantile regression NN, to address error propagation and accumulation in multi-step flood probability density forecasts. Additionally, a hybrid error correction model [265] is proposed to enhance streamflow forecast accuracy. The hybrid model integrates the HBV model with the ML algorithm through a data assimilation technique.
In general, integrating physics knowledge into ML for uncertainty estimation has the potential to better characterize uncertainty. For instance, ML surrogate models might yield physically inconsistent predictions; integrating physical principles can help alleviate this problem. Additionally, ML can reduce the uncertainty of rainfall-runoff models by effectively extracting features and identifying significant variables [274]. The reduced data needs of ML due to constraints for adherence to known rainfall-runoff processes could alleviate some of the computational costs associated with NNs. However, rainfall-runoff modeling remains a complex process, and defining a simple, unified benchmark—encompassing effective and publicly available datasets, robust metrics, and comprehensive baseline tools—for different PaML methods is a significant challenge [262]. Furthermore, current PaML-based methods mainly focus on PDgML and PaHL, which face substantial challenges in addressing effectiveness, computational complexity, and usability in uncertainty estimation for rainfall-runoff forecasts. Therefore, deepening the integration of physical information from the rainfall-runoff process into ML frameworks or optimization processes (such as PeML and PiML) to achieve low uncertainty, end-to-end, and effective rainfall-runoff process modeling will be a crucial direction for future research.